CHEM 1405-Experiment 8: Chemical Reactions Section: Date: Name: Post-lab Questions Write balanced equations for each of the following reactions: gle replacement reaction between aluminum metal and aqueous nickel(II) chloride. (b) Single replacement reaction between calcium metal and hydrochloric acid. ⓒ Duble replacement and decomposition reaction between sodium, bicarbonate and acetic acid (Hin: double replacement firest, then carbonic acid decomposes into CO; and another product - 3 products in total 2 In each of the following changes, indicate whether the starting substance is getting oxidized or reduced (assume that another, unnamed substance accounts for the oxidizing or reducing agent). (a) Formaldehyde, HCHO is converted to formic acid, HCOOH. (b) Acetone, CH COC Hs is converted to iso-propyl alcohol, CH:CHOHCHs Suppose the mass of one mole of a substance is 45.0 g. How many moles of the substance is present in 0.568 g of the sample? 3 4 How many molecules of water are in 2.55 moles of water? 5 Challenge: How many molecules of carbon dioxide are in 4.56 g of carbon dioxide?

media%2F944%2F944c5541-25bc-4c1b-a30a-4b
Answer #1

Ans1. (a) 2Al(s) + 3 NiCl2(aq) = 2AlCl3(aq) + 3 Ni (s) ............replacement reaction

(b) Ca(s) + 2 HCl(aq) = CaCl2 (aq) + H2

(c) Na2CO3(aq) + 2CH3COOH(aq) = 2CH3COONa + CO2 + H2O

Ans 2.(a) starting material HCHO is getting oxidized and converted to formic acid

(b) starting material acetone is getting reduced to form isopropyl alcohol.

Ans3 45g of substance = 1 mol

0.568g of sample contains = 1 * 0.568 / 45 = 0.0126 mol

Ans 4. 1 mole of H2O contains 6.022 * 10^23 molecules

and hence 2.55 moles of H2O contains 6.022 * 2.55 * 10^23 = 1.535 * 10^24 molecules of H2O